# CASCRNet: An Atrous Spatial Pyramid Pooling and Shared Channel Residual based Network for Capsule Endoscopy

### v1

@@ 1 Introduction
As proposed by authors of [3], the challenge in hand invites contestants to create a model capable of automatically classifying abnormalities captured in video capsule endoscopy (VCE) video frames. The dataset [4] contains 37607 training images and 16132 validation images belonging to 10 classes each viz. Angioectasia, Bleeding, Erosion, Erythema, Foreign Body, Lymphangiectasia, Normal, Polyp, Ulcer and Worms. The testing set consists of 4385 images. The training and validation sets are highly imbalanced with Normal class being the large chunk of the data.
The authors of [3] have conducted a preliminary analysis utilizing a variety of machine-learning and deep-learning models such as ResNet50 [5], VGG16 [6], Support Vector Machines (SVM) [7] among others. DeepLab [8] introduced atrous convolutions, also known as dilated convolutions, which increase the receptive field of the network without requiring additional parameters or downsampling, making them effective for dense prediction tasks like semantic segmentation. A key component of the proposed model is the incorporation of atrous convolution. RCCGNet [9] utilized the concept of shared channel residual blocks for the classification of renal cell carcinoma and achieved good results while keeping the model compact. This study uses this approach in the proposed model with a few additions and modifications. This GitHub [^1] https://github.com/Manvith-Prabhu/Capsule-Vision-2024.git repository contains the source code for this research.
@@ 2.3 Proposed CASCRNet
The proposed model, CASCRNet, leverages the same shared channel residual block concept depicted in Figure 2. An Atrous Spatial Pyramid Pooling block was incorporated to further enhance the model’s ability to understand images. The LeakyReLU [12] activation function with an alpha value of 0.01 was utilized. In addition to this, the Adam optimizer [13] was used, whose learning rate started with a value of 0.001 and was halved each time it reached a plateau. Dilated convolutions were applied within the residual blocks. Additionally, focal loss was implemented to improve performance. The architecture of CASCRNet is illustrated in Figure 3.
@@ 6 Acknowledgments
As participants in the Capsule Vision 2024 Challenge, we fully comply with the competition’s rules as outlined in [3]. Our AI model development is based exclusively on the datasets provided in the official release in [4].

### v2

@@ 1 Introduction
As proposed by authors of [3], the challenge in hand invites contestants to create a model capable of automatically classifying abnormalities captured in video capsule endoscopy (VCE) video frames. The dataset [4] contains 37607 training images and 16132 validation images belonging to 10 classes each viz. Angioectasia, Bleeding, Erosion, Erythema, Foreign Body, Lymphangiectasia, Normal, Polyp, Ulcer and Worms. The testing set [5] consists of 4385 images. The training and validation sets are highly imbalanced with Normal class being the large chunk of the data.
The authors of [3] have conducted a preliminary analysis utilizing a variety of machine-learning and deep-learning models such as ResNet50 [6], VGG16 [7], Support Vector Machines (SVM) [8] among others. DeepLab [9] introduced atrous convolutions, also known as dilated convolutions, which increase the receptive field of the network without requiring additional parameters or downsampling, making them effective for dense prediction tasks like semantic segmentation. A key component of the proposed model is the incorporation of atrous convolution. RCCGNet [10] utilized the concept of shared channel residual blocks for the classification of renal cell carcinoma and achieved good results while keeping the model compact. This study uses this approach in the proposed model with a few additions and modifications. This GitHub [^1] https://github.com/Manvith-Prabhu/CASCRNet.git repository contains the source code for this research. We were ranked #2 on Validation set and #15 on testing set performance.
@@ 2.3 Proposed CASCRNet
The proposed model, CASCRNet, leverages the same shared channel residual block concept depicted in Figure 2. An Atrous Spatial Pyramid Pooling block was incorporated to further enhance the model’s ability to understand images. The LeakyReLU [13] activation function with an alpha value of 0.01 was utilized. In addition to this, the Adam optimizer [14] was used, whose learning rate started with a value of 0.001 and was halved each time it reached a plateau. Dilated convolutions were applied within the residual blocks. Additionally, focal loss was implemented to tackle the class imbalance issue. The architecture of CASCRNet is illustrated in Figure 3.
@@ 6 Acknowledgments
As participants in the Capsule Vision 2024 Challenge, we fully comply with the competition’s rules as outlined in [3]. Our AI model development is based exclusively on the datasets provided in the official release in [4] and [5].